OpenStudy (anonymous):

(1/6,-1/3) and (5/6, 5) what is the equation of the line Y=?

OpenStudy (anonymous):

\[\frac{y-y1}{y2-y1} = \frac{x-x1}{x2-x1}\] So you label your points (1/6,-1/3) as (x1,y1) and (5/6, 5) as (x2,y2) and feed them in to the above formula and simplify, you will get your equation.